I don't understand this question. How do you solve it? simplify f(x) = the quantity 2 x-squared plus 3x plus 1 all over 2x plus 1. Using complete sentences, explain why f(1) = 2, f(0) = 1, and f(-1) = 0, yet f(-one¨Chalf) is undefined. Make sure to show your work.
Answer by ad_alta(240)

Let's rewrite the function f(x)=(2x^2+3x+1)/(2x+1), since we can factor 2x^2+3x+1 into (2x+1)(x+1). We can always cancel the factor of 2x+1, except at -1/2 where the function is undefined (since we can't divide by zero). Thus, f(1)=1+1=2, f(0)=0+1=1, and f(-1)=-1+1=0.